# Hyperelliptic Shimura curves and $L$-functions of central vanishing order at least 3

### v1

@@ 1. Introduction
Let $F={\mathbb{Q}}\left(\cos\frac{1}{16}\pi\right)$. Let $\mathfrak{p}$} be the unique prime of $F$ above 2. Let $A_{1},A_{2},A_{3}$ be 4-dimensional simple modular abelian varieties of ${\mathrm{GL}}_{2}$-type over $F$ and of conductor $\mathfrak{p}$}. Then
@@
Indeed, there are exactly four modular abelian varieties of ${\mathrm{GL}}_{2}$-type over $F$ of conductor $\mathfrak{p}$} and dimension 4, which are all simple. Each of them correspond to a Galois orbit of size 4 of Hilbert newforms over $F$. Then the fact that the triple product $L$-function has a holomorphic continuation to $s\in{\mathbb{C}}$ follows from [5, 17].
We actually prove a slightly more general version of 1.1 in terms of Hilbert newforms over $F$. In 7.2, we show that for any 3 of the 16 Hilbert newforms (in the four Galois orbits of size 4) above, their triple product $L$-function has central vanishing order at least 3.
The proof is similar to the one of Gross and Zagier in spirit. They used their celebrated formula on modular curves, which relates $L$-functions of modular forms to the heights of (Heegner) CM 0-cycles. We use an analog proved by Yuan, S. Zhang, and W. Zhang [24] for some Shimura curves, which relates triples product $L$-functions to the heights of modified diagonal cycles. Both of us show that our cycles accidently vanish for some reasons.
Besides, the specific modular curve used for Gross and Zagier’s order 3 example is $X_{0}(37)$ (see [26]) and it has an exceptional hyperelliptic involution found by Ogg [16]. So does the Shimura curve used by us, which is the work of Dembélé [4]. It might be a coincidence, since the hyperellipticity is essential for our result, but it is not used in [11, 26]. However, in 4.7, we show that hyperellipticity can be used to show vanishing of CM 0-cycles. And in 6.2, we show that hyperellipticity is also related to higher central vanishing of the $L$-functions of single Hilbert newforms.
Finally, we want to remark that there are only finitely many hyperelliptic Shimura curve, as shown in [20]. This is also a facet of the expectation that $L$-functions of central vanishing order at least 3 are very rare.
@@ Acknowledgments
The author would like to thank Peter Sarnak for raising the question of on possible analogs of Gross–Zagier’s example in higher dimensions. The author is also grateful to Yifeng Liu and Wei Zhang for many helpful discussions.
@@ 2. Triple product of abelian motives
Let $A$ be an abelian variety over a field $F$ of dimension $g$, with dual abelian variety $A^{\vee}$. Let ${\mathfrak{pr}}\in{\mathrm{Ch}}_{g}(A\times A)$ be the (2g1)(2g-1)-th Deninger–Murre projector, that is denoted by $\pi_{2g-1}$ in [3, Theorem 3.1]. For a variety $X$ over $F$, the abelian group structure of $A$ induces an abelian group structure on ${\mathrm{Mor}}(X,A)$, the set of morphisms of $F$-varieties. Then by regarding $X\times A$ as an abelian scheme over $X$, [23, Corollary A.3.2] implies that the following map is a homomorphism of abelian groups:
@@
For an idempotent $\iota\in{\mathrm{End}}(A)_{\mathbb{C}}$, by [3, Proposition 3.3], ${\mathfrak{pr}}_{\iota}$ is a projector. Moreover, if $\iota$ is in the center of ${\mathrm{End}}(A)_{\mathbb{C}}$, one can easily check that its image $\iota^{\dagger}$ under a Rosati involution on ${\mathrm{End}}(A)_{\mathbb{C}}$ associated with a polarization does not depend on the choice of the polarization.
@@
The proof is the same as [21, Corollary 2.2.3]. ∎
@@
The proof is the same as [21, Corollary 2.2.7 (1)]. ∎
@@
The relation between the above two propositions and [21, Corollary 2.2.3 , Corollary 2.2.7 (1)] is not just similar. Indeed, if $A$ is simple, we have an isomoprhism between the motives $(A,{\mathfrak{pr}}_{\iota})$ and $(X,\delta_{f})$ defined in [21, Section 2.4][1][1][1]In loc. cit., we used $\phi$ instead of $f$ to denote a morphism from $X$ to $A$. . This isomoprhism can be easily proved using the decomposition of $\delta_{f}$ in (the summand on the right hand side of) [23, (2.13)] and the nilpotency of Kimura [13, Proposition 7.5].
@@ 3. Shimura curves
where $L_{U}^{\circ}$ is the restriction of $L_{U}$ to any geometrically connected component of $X_{U}$. In [23, Remark 3.2.4], we further noticed that if $X_{U}$ is a modular curve, then ${\mathfrak{c}}_{U}$ is a multiple of $\lambda_{U}$. So
@@
Finally in this section, we explain the relation between the divisors defined above. Below we simplify notations and let $\lambda,{\mathfrak{c}},{\mathfrak{e}}$ denote $\lambda_{U},{\mathfrak{c}}_{U},{\mathfrak{e}}_{U}$ if $U$ is clear from the context.
For $z\in{\mathrm{Ch}}_{1}(X_{U})$, by [23, Corollary 2.5.3], we have
[…]
We need another class close to $\lambda$ to discuss (modified) diagonals. Assume $\deg(K_{U})>0$. Let
@@
If $\deg{\mathfrak{e}}>0$, then by (3.2),
(3.4) $\xi=\lambda\text{ if and only if }[{\mathfrak{e}}]=(\deg{\mathfrak{e}})\lambda.$
@@ 4. Diagonals and hyperellipticity
where the transition maps in the inverse limit are pushforwrds by the natural morphisms between Shimura curves. Let $B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$ act on $\overline{{\mathrm{Ch}}}(n)$ (from left) by pullback. Let $\pi_{i}$ be of the form $\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2), which is a representation of $B^{\times}({\mathbb{A}}_{\mathrm{f}})$.
Theorem 4.1 ([23, Theorem 3.3.2]).
Let $H\subset B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$ be a subgroup such that $\boxtimes_{i=1}^{n}\pi_{i}$ has no nonzero $H$-invariant linear forms. Then for $z=(z_{U})_{U}\in{\mathrm{Ch}}(n)$ invariant by $H$, we have
$\left(\mathop{\otimes}_{i=1}^{n}{\mathfrak{pr}}_{f_{i}}\right)_{*}z=0$
for all $U$ and all $f_{i}\in\pi_{i}^{U}$.
@@
$\Delta_{n}]_{f_{1},f_{2},...,f_{n}}=\left({\mathfrak{pr}}_{f_{1}}\otimes{\mathfrak{pr}}_{f_{2}}\otimes...\otimes{\mathfrak{pr}}_{f_{n}}\right)_{*}[\Delta_{n}]$
@@
Corollary 4.2 ([23, Corollary 3.3.4]).
Assume that $\boxtimes_{i=1}^{n}\pi_{i}$ has no nonzero diagonal-$B^{\times}({\mathbb{A}}_{\mathrm{f}})$-invariant linear forms, then
$\Delta_{n}]_{f_{1},f_{2},...,f_{n}}=0$
for all $U$ and all $f_{i}\in\pi_{i}^{U}$.
Here are some easy examples for applying this corollary.
@@
(1) If $n=1$, let $H=B^{\times}({\mathbb{A}}_{\mathrm{f}})$ and we have $\Delta_{1}]_{f_{1}}=0$ for all $f_{1}\in\pi_{1}^{U}$.
(2) If $n=2$, $H$ is the diagonal $B^{\times}({\mathbb{A}}_{\mathrm{f}})$, and $z_{U}=[\Delta_{2}]$, then $\Delta_{2}]_{f_{1},f_{2}}=0$ if $\pi_{1}\not\simeq\pi_{2}^{\vee}$.
(3) If the product of the central characters of $\pi_{i}$’s is not trivial, then $\Delta_{n}]_{f_{1},f_{2},...,f_{n}}=0$.
@@
Assume every geometrically connected component of $X_{U}$ is hyperelliptic. Then
$\Delta_{3}]_{f_{1},f_{2},f_{3}}=0$
@@
We prove it in the end of this section. We need some preparations.
Lemma 4.5.
$\Delta_{3}]_{f_{1},f_{2},f_{3}}=-\sum_{\begin{subarray}{c}I\coprod J=\{1,2,3\}\\ I=\{a,b\},J=\{c\}\end{subarray}}p_{I}^{*}[\Delta_{2}]_{f_{a},f_{b}}\cdot p_{J}^{*}{\mathfrak{pr}}_{f_{c},*}\xi$
@@
By hyperellipticity and [10, 4.5], $\Delta_{\xi}]=0$. The proposition follows from 4.3 (1)(2). ∎
By 2.1 and 3.2 (2), we have the following lemma.
Lemma 4.6.
If $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}$ and $f_{1}=f_{2}=f_{3}=:f$, then $\Delta_{3}]_{f,f,f}=0$.
Using the last two lemmas, we are able prove the following.
@@
Assume every geometrically connected component of $X_{U}$ is hyperelliptic. Then $\xi=\lambda$, ${\mathfrak{e}}]=(\deg{\mathfrak{e}})\lambda$, ${\mathfrak{pr}}_{f,*}\xi=0$ and ${\mathfrak{pr}}_{f,*}[{\mathfrak{e}}]=0$ for every $\pi=\pi_{A,\iota}$ and every $f\in\pi^{U}$.
Proof.
By (3.3) and (3.4), the four equations in the corollary are equivalent to each other. We prove the last two of them. By (3.5), the last two are equivalent to each other for each $f$.
If $\pi$ is not self-dual so that $\pi$ has nontrivial central character, then ${\mathfrak{pr}}_{f,*}{\mathfrak{e}}=0$ by using the formulation in [25, 1.3.2] with invariant linear forms on $\pi$. So we assume that $\pi$ is self-dual and want to prove that ${\mathfrak{pr}}_{f,*}\xi=0$ instead.
We use 4.5 (1) with with $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=\pi$ and $f_{1}=f_{2}=f_{3}=:f$. Indeed, by 2.1, the left hand side of the equation in (3) of the lemma is 0. Now we analyze the terms on the right hand side, and one of them is $p_{1,2}^{*}[\Delta_{2}]_{f,f}\cdot p_{3}^{*}{\mathfrak{pr}}_{f,*}\xi,$ which is just $\Delta_{2}]_{f,f}\times{\mathfrak{pr}}_{f,*}\xi$.
Claim: there is a cycle $z$ in ${\mathrm{Ch}}_{2\dim A-1}(A\times A)$ whose intersection number with $\Delta_{2}]_{f,f}$ is nonzero. Since the cohomology class of $\xi$ is 0, the intersection of the right hand side of the equation in 4.5 (1) with $z\times[A]$ is
$([\Delta_{2}]_{f,f}\times{\mathfrak{pr}}_{f,*}\xi)\cdot(z\times[A]).$
Then the push forward by $p_{3}$ of this Chow cycle is a multiple of ${\mathfrak{pr}}_{f,*}\xi$. But it is 0 by left hand side being 0.
Now we prove the claim. Since $\pi$ is self dual, by pulling back to $C$ (as in the summand on the right hand side of [23, (2.13)]), we see that the Betti cohomology class of $\Delta_{2}]_{f,f}$ in $H^{1}(A_{\mathbb{C}})\otimes H^{1}(A_{\mathbb{C}})$, which actually is in $\iota_{*}H^{1}(A_{\mathbb{C}})\otimes\iota_{*}H^{1}(A_{\mathbb{C}})$,) is nonzero. By the nonvanishing of the intersection pairing between $H^{1}(A_{\mathbb{C}})$ and $H^{2\dim A-1}(A_{\mathbb{C}})$, and thus their tensor squares, we can find $z$ as in the claim. ∎
@@
Note that elliptic points are CM points and $\mathfrak{e}$} is a sum of CM cycles. Then ${\mathfrak{pr}}_{f,*}{\mathfrak{e}}=0$ gives an explicit linear relation between the images by ${\mathfrak{pr}}_{f,*}$ of the CM zero cycles on $A$. This is trivial if the images are all 0, which is the case if $L^{\prime}(1/2,\pi)=0$ by the general Gross–Zagier formula [11, 25]. (In the corollary, our $\pi$ denotes the finite part of an automorphic representation of $B^{\times}$. In this remark, $L(s,\pi)$ denotes the $L$-function of this automorphic representation of $B^{\times}$.) So we assume $L^{\prime}(1/2,\pi)\neq 0$. For example, let $X_{U}$ be the classical modular curve $X_{0}(37)$, which is famously hyperelliptic by Ogg [16]. There is a unique such $\pi$, corresponding to the newform $f$ with LMFDB label 37.2.a.a. There are two elliptic points of order 2 (resp. 3) on $X_{U}$, and let ${\mathfrak{e}}_{2}$ (resp. ${\mathfrak{e}}_{3}$) be their sum. Then ${\mathfrak{pr}}_{f,*}{\mathfrak{e}}=0$ gives
(4.2) $\frac{1}{2}{\mathfrak{pr}}_{f,*}{\mathfrak{e}}_{2}+\frac{2}{3}{\mathfrak{pr}}_{f,*}{\mathfrak{e}}_{3}=0.$
This could also be read from the modular form of weight $3/2$ corresponding to $f$, by the Gross–Kohnen–Zagier formula [8]. We leave it to the interested reader. Besides, (4.2) and the version of the Gross–Zagier formula [1, Theorem 1.1] implies that the finite part of the quadratic twist $L$-functions of $\pi$ satisfies
$\frac{2^{2}\cdot\sqrt{4\cdot 2^{2}}\cdot L^{\infty}(1/2,\pi\otimes\chi_{4})}{3^{2}\cdot\sqrt{3}\cdot L^{\infty}(1/2,\pi\otimes\chi_{3})}=\frac{16}{9}.$
One can also check that this is true up to very high precision using $L$-values read from LMFDB [14].
Proof of $4.$
It follows from 4.5 and 4.7. ∎

### v2

@@ 1. Introduction
Let $F={\mathbb{Q}}(\zeta_{32}+\zeta_{32}^{-1})$, the maximal totally real subfield of the cyclotomic field of 32nd roots of unity. Let $\mathfrak{p}$} be the unique prime of $F$ above 2. Let $A_{1},A_{2},A_{3}$ be pairwise non-isomorphic 4-dimensional simple modular abelian varieties of ${\mathrm{GL}}_{2}$-type over $F$ and of conductor $\mathfrak{p}$}. Then
@@
Indeed, there are exactly four modular abelian varieties of ${\mathrm{GL}}_{2}$-type over $F$ of conductor $\mathfrak{p}$} and dimension 4, which are all simple. They exactly correspond to the only 4 Galois orbits of Hilbert newforms over $F$ with Atkin–Lehner sign 1-1. And each Galois orbit is of size 4. Then the fact that the triple product $L$-function has a holomorphic continuation to $s\in{\mathbb{C}}$ follows from [4, 16].
We actually prove a slightly more general version of 1.1 in terms of Hilbert newforms over $F$. In 8.2, we show that for any 3 pairwise distinct newforms of the 16 Hilbert newforms (in the four Galois orbits of size 4) above, their triple product $L$-function has central vanishing order at least 3.
In general, when some of the 3 newforms are the same, the triple product $L$-function is decomposable. See Section 6 . In this case, using usual modular forms over $\mathbb{Q}$}, we get many degree 6 $L$-functions with central vanishing order at least 2 (7.2).
Our proof of 1.1 is similar to the one of Gross and Zagier in spirit. They used their celebrated formula on modular curves, which relates $L$-functions of modular forms to the heights of (Heegner) CM 0-cycles. We use an analog proved by Yuan, S. Zhang, and W. Zhang [24] for some Shimura curves, which relates triples product $L$-functions to the heights of modified diagonal cycles on the triple product of Shimura curves. Both of us show that our cycles accidentally vanish for some reasons.
In our case, our cycles vanish due to an exceptional hyperelliptic involution, which is the work of Dembélé [3]. See Section 8. Here, an involution is exceptional if it does not come from an Atkin–Lehner involution. (In general, if an Atkin–Lehner involution is a hyperelliptic involution, it is not useful to produce higher central vanishing of $L$-functions. See 5.8.) Interestingly, the specific modular curve used for Gross and Zagier’s order 3 example, which is $X_{0}(37)$, is of genus 2. And its hyperelliptic involution is exceptional. It might be a coincidence, since the hyperellipticity is essential for our result, but it is not used in [10, 26].
Finally, we want to remark that there are only finitely many hyperelliptic Shimura curve, as shown in [19], let along the ones satisfying our specific criterions Section 5. Moreover, we are mostly interested in the ones of genus at least 3 due to the decomposability of triple product $L$-functions mentioned above. They are even rarer. For example, there are no such modular curves as we will see in Section 7. Such rareness is also a facet of the expectation that $L$-functions of central vanishing order at least 3 are very rare.
Structure of the paper
In Section 2 and Section 3, we review some background on abelian varieties and Shimura curves, respectively. In Section 4, we relate modified diagonals to the hyperellipticity of Shimura curves. In Section 5, we give criteria for triple product $L$-functions to have central vanishing order at least 3. In Section 6, we consider the cases where the triple product $L$-function is decomposable. In Section 7, we study examples of hyperelliptic modular curves and obtain many degree-6 $L$-functions over $\mathbb{Q}$} with central vanishing order at least 2. Finally, in Section 8, we study the hyperelliptic Shimura curve of Dembélé and prove 1.1.
@@ Acknowledgments
The author would like to thank Peter Sarnak for raising the question of possible analogs of Gross–Zagier’s example in higher dimensions during a talk by the author at Princeton. The author is grateful to Yifeng Liu and Wei Zhang for many helpful discussions in the early stages of this work. The author also thanks Benedict Gross and Adam Logan for valuable communications that helped correct a mistake in the first version of this paper.
@@ 2. Triple product of abelian motives
Let $A$ be an abelian variety over a field $F$ of dimension $g$, with dual abelian variety $A^{\vee}$. Let ${\mathfrak{pr}}\in{\mathrm{Ch}}_{g}(A\times A)$ be the (2g1)(2g-1)-th Deninger–Murre projector, that is denoted by $\pi_{2g-1}$ in [2, Theorem 3.1]. For a variety $X$ over $F$, the abelian group structure of $A$ induces an abelian group structure on ${\mathrm{Mor}}(X,A)$, the set of morphisms of $F$-varieties. Then by regarding $X\times A$ as an abelian scheme over $X$, [22, Corollary A.3.2] implies that the following map is a homomorphism of abelian groups:
@@
For an idempotent $\iota\in{\mathrm{End}}(A)_{\mathbb{C}}$, by [2, Proposition 3.3], ${\mathfrak{pr}}_{\iota}$ is a projector. Moreover, if $\iota$ is in the center of ${\mathrm{End}}(A)_{\mathbb{C}}$, one can easily check that its image $\iota^{\dagger}$ under a Rosati involution on ${\mathrm{End}}(A)_{\mathbb{C}}$ associated with a polarization does not depend on the choice of the polarization.
@@
The proof is the same as [20, Corollary 2.2.3]. ∎
@@
The proof is the same as [20, Corollary 2.2.7 (1)]. ∎
@@
The relation between the above two propositions and [20, Corollary 2.2.3 , Corollary 2.2.7 (1)] is not just similar. Indeed, if $A$ is simple, we have an isomoprhism between the motives $(A,{\mathfrak{pr}}_{\iota})$ and $(X,\delta_{f})$ defined in [20, Section 2.4][1][1][1]In loc. cit., we used $\phi$ instead of $f$ to denote a morphism from $X$ to $A$. . This isomoprhism can be easily proved using the decomposition of $\delta_{f}$ in (the summand on the right hand side of) [22, (2.13)] and the nilpotency of Kimura [13, Proposition 7.5].
@@ 3. Shimura curves
where $L_{U}^{\circ}$ is the restriction of $L_{U}$ to any geometrically connected component of $X_{U}$. In [22, Remark 3.2.4], we further noticed that if $X_{U}$ is a modular curve, then ${\mathfrak{c}}_{U}$ is a multiple of $\lambda_{U}$. So
@@
Below we simplify notations and let $\lambda,{\mathfrak{c}},{\mathfrak{e}}$ denote $\lambda_{U},{\mathfrak{c}}_{U},{\mathfrak{e}}_{U}$ if $U$ is clear from the context. Assume $\deg(K_{U})>0$. Let
@@ 4. Diagonals and hyperellipticity
where the transition maps in the inverse limit are pushforwrds by the natural morphisms between Shimura curves. Let $B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$ act on $\overline{{\mathrm{Ch}}}(n)$ (from left) by pullback. Let $\pi_{i}$ be of the form $\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2), which is a representation of $B^{\times}({\mathbb{A}}_{\mathrm{f}})$. We have a linear map $\pi_{i}^{U}\to{\mathrm{Ch}}_{1}(X_{U}\times A_{i})$, sending $f_{i}$ to ${\mathfrak{pr}}_{f_{i}}$. Let
$\boxtimes_{i=1}^{n}\pi_{i}^{U}$ $\to\otimes_{i=1}^{n}{\mathrm{Ch}}_{1}(X_{U}\times A_{i}),$
$f$ $\mapsto{\mathfrak{pr}}_{f}.$
be the tensor product of these maps.
The following is a slight generalization of [22, Theorem 3.3.2]. The proof is the same.
Theorem 4.1.
Let $H\subset B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$ be a subgroup. Let $\Pi\subset\boxtimes_{i=1}^{n}\pi_{i}$ be a subspace be stable by $H$. If $\Pi$ has no nonzero $H$-invariant linear forms, then for $z=(z_{U})_{U}\in{\mathrm{Ch}}(n)$ invariant by $H$, we have
${\mathfrak{pr}}_{f,*}z=0$
for all $f\in\Pi$.
In particular, the theorem applies to a sub-$B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$-representation $\Pi$ of $\boxtimes_{i=1}^{n}\pi_{i}$.
@@
$\Delta_{n}]_{f}={\mathfrak{pr}}_{f,*}[\Delta_{n}]$
@@
Accordingly, we have the following slight generalization of [22, Corollary 3.3.4].
Corollary 4.2.
If a sub-$B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$-representation $\Pi$ of $\boxtimes_{i=1}^{n}\pi_{i}$. has no nonzero diagonal-$B^{\times}({\mathbb{A}}_{\mathrm{f}})$-invariant linear forms, then
$\Delta_{n}]_{f}=0$
for all $f\in\Pi$.
Remark 4.3.
In the case $n=3$ and $\Pi=\boxtimes_{i=1}^{n}\pi_{i}$, this condition has been extensively studied in [19].
Here are some easy examples for applying this corollary to $\Pi=\boxtimes_{i=1}^{n}\pi_{i}$.
@@
(1) If $n=1$, let $H=B^{\times}({\mathbb{A}}_{\mathrm{f}})$ and we have $\Delta_{1}]_{f_{1}}=0$.
(2) If $n=2$, $H$ is the diagonal $B^{\times}({\mathbb{A}}_{\mathrm{f}})$, and $z_{U}=[\Delta_{2}]$, then $\Delta_{2}]_{f_{1}\otimes f_{2}}=0$ if $\pi_{1}\not\simeq\pi_{2}^{\vee}$.
(3) If the product of the central characters of $\pi_{i}$’s is not trivial, then $\Delta_{n}]_{f_{1}\otimes f_{2},...\otimes f_{n}}=0$.
@@
$\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=-\sum_{\begin{subarray}{c}I\coprod J=\{1,2,3\}\\ I=\{a,b\},J=\{c\}\end{subarray}}p_{I}^{*}[\Delta_{2}]_{f_{a}\otimes f_{b}}\cdot p_{J}^{*}{\mathfrak{pr}}_{f_{c},*}\xi$
@@
(2) Further assume $\pi_{3}\not\simeq\pi_{1}^{\vee}$ and $\pi_{3}\not\simeq\pi_{2}^{\vee}$. Then
$\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=-p_{1,2}^{*}[\Delta_{2}]_{f_{1}\otimes f_{2}}\cdot p_{3}^{*}{\mathfrak{pr}}_{f_{3},*}\xi,$
@@
By hyperellipticity and [9, 4.5], $\Delta_{\xi}]=0$. The proposition follows from 4.4 (1)(2). ∎
Combining (1) and (2) of the proposition with (3.3), we have a corollary.
@@
(1) Assume every geometrically connected component of $X_{U}$ is hyperelliptic. Then $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=0$ for $f_{i}\in\pi_{i}^{U}$ such that ${\mathfrak{pr}}_{f_{i},*}[{\mathfrak{e}}]=0$.
(2) Further assume $\pi_{3}\not\simeq\pi_{1}^{\vee}$ and $\pi_{3}\not\simeq\pi_{2}^{\vee}$. Then $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=0$ for $f_{i}\in\pi_{i}^{U}$ such that ${\mathfrak{pr}}_{f_{3},*}[{\mathfrak{e}}]=0.$
@@
Note that elliptic points are CM points and $\mathfrak{e}$} is a sum of CM cycles. Then the equation ${\mathfrak{pr}}_{f,*}[{\mathfrak{e}}]=0$, if holds, can usually be verified by local conditions. See Section 7.